A 10 ft3 tank contains gas at a pressure of 500 psia, temperature of 85 0F and

a mass of 25 lbm. A part of the gas was discharged and the temperature and

pressure changed to 70 0F and 300 psia, respectively. Heat was applied and

the temperature was back to 85 0F. Find the final mass, volume and pressure

of the gas. (15.43 lbm, 10 ft3, 308.5 psia)


1
Expert's answer
2022-02-14T14:25:01-0500

Gives

"T_3=85°F\\T_2=70°F"

We know that

"PV=mRT"

"\frac{m_1RT_1}{P_1}=\frac{m_2RT_2}{P_2}"

"\frac{P_2}{T_2}=\frac{P_3}{T_3}"

"P_3=\frac{T_3}{T_2}\times P_2"


"P_3=(\frac{85+460}{70+460})\times300=308.5Psia"

"\frac{m_1}{P_1}=\frac{m_3}{P_3}"


"m_3=\frac{P_3}{P_1}\times m_1=\frac{25\times308.5}{500}=15.43lbm"

Now volume

"V_1=V_2=V_3=10ft^3"